### v1

# Weighted blow-up desingularization of plane curves

Abstract.
Given a singular hypersurface in a regular 2-dimensional scheme essentially of finite type over a field, we construct an embedded resolution of singularities by tame Artin stacks. This applies in particular to plane curves. The key tool is Hironaka’s characteristic polyhedron from which we determine a suitable center for a stack-theoretic weighted blow-up. In particular, we show that the order strictly decreases at every point lying above the center. In contrast to earlier work by the authors, we do not have to restrict to perfect base fields and we deduce an inductive argument, despite the fact that higher dimensional tangent spaces arise, by taking torus actions and equivariant centers into account. Hence, we work with weighted blow-ups of tame Artin stacks and we do not have to pass to multi-weighted blow-ups of Deligne–Mumford stacks.
Key words and phrases:
Resolution of singularities, Artin stacks, curve singularities, weighted blow-ups.
2020 Mathematics Subject Classification:
14H20, 14B05, 14M25, 14D23
1. Introduction
This paper is a continuation of [AQS], where we revisited the resolution of singularities of plane curves using stack-theoretic blow-ups. In that paper we considered a singular curve $C\subset S$ embedded in a smooth surface over a field $K$ which we assumed perfect. We showed in [AQS, Theorem 1.1] that there is a natural weighted blow-up $S^{\prime}\to S$, where the proper transform $C^{\prime}\subset S^{\prime}$ of $C$ has lower order. The surface $S^{\prime}$ is, however, in general an Artin stack of the form $[\mathcal{W}^{\prime}/\mathbb{G}_{m}]$, where the action has possibly inseparable stabilizers, and $\mathcal{W}^{\prime}$ is a 3-dimensional scheme. In order to be able to use induction, in [AQS, Theorem 1.2] we replaced $S^{\prime}$ by another stack-theoretic blowup $S^{\prime\prime}\to S$, which is a Deligne–Mumford stack, allowing us to repeat the process.
In this paper we take a more direct approach, using the weighted blowup $S^{\prime}=[\mathcal{W}^{\prime}/\mathbb{G}_{m}]$ itself. In order to be able to repeat blowing up $S^{\prime}$, we expand the generality where we allow higher dimensional schemes, but require them to possess a torus actions. The precise setup is as follows:
Theorem 1.1.
Let $\mathcal{Z}\subset\mathcal{W}$ be a hypersurface in a regular scheme $\mathcal{W}$ of dimension $n+2$ which is essentially of finite type over an arbitrary field $k$. Assume given an action of $T=\mathbb{G}_{m}^{n}$ on $\mathcal{W}$ stabilizing $\mathcal{Z}$, having finite, possibly non-reduced, stabilizers. Let $\mathcal{Q}\subset\mathcal{Z}$ be a singular closed orbit such that the reduction $\mathcal{Z}_{\rm red}$ of $\mathcal{Z}$ is singular along $\mathcal{Q}$.
  1. (1)
    There is a well-defined, unique $T$-invariant center $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ of maximal invariant $(a_{1},a_{2})$ that is admissible for $\mathcal{Z}\subset\mathcal{W}$ and supported along $\mathcal{Q}$, where $a_{1}$ is the order of $\mathcal{Z}$ along $\mathcal{Q}$. Moreover, $J$ is stable under base change to separable field extensions of $k$.
  2. (2)
    Let $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$ be the associated reduced center. Let $B\to\mathcal{W}$ be the associated degeneration to the normal cone of $\bar{J}$, with a natural open subscheme $\mathcal{W}^{\prime}:=B_{+}\subset B$ admitting a $T^{\prime}=T\times\mathbb{G}_{m}$-action lifting the $T$-action on $\mathcal{W}$. Then the action of $T^{\prime}$ on $\mathcal{W}^{\prime}$ has finite stabilizers, and the morphism $[\mathcal{W}^{\prime}/T^{\prime}]\to[\mathcal{W}/T]$ is the weighted blow-up along $\bar{J}$, which is a proper and birational morphism of tame Artin stacks.
  3. (3)
    Let $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}$ be the proper transform of $\mathcal{Z}$. Then the order of $\mathcal{Z}^{\prime}$ at every point lying over $\mathcal{Q}$ is strictly smaller than $a_{1}$. In particular, the invariant defined in Part (1) of the center associated to $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}$ at every point of $\mathcal{W}^{\prime}$ lying over $\mathcal{Q}$ is strictly smaller than $(a_{1},a_{2})$.
Iterating $\mathcal{W}^{\prime}\to\mathcal{W}$, one obtains in finitely many steps $\mathcal{W}^{(m)}\to\mathcal{W}^{(m-1)}$, with the reduction $\mathcal{Z}^{(m)}_{\operatorname{red}}\subset\mathcal{W}^{(m)}$ nonsingular. Taking quotients $[\mathcal{Z}^{(m)}/T^{(m)}]_{\operatorname{red}}\subset[\mathcal{W}^{(m)}/T^{(m)}]$ we obtain an embedded resolution of singularities of $[\mathcal{Z}/T]\subset[\mathcal{W}/T]$ by tame Artin stacks.
Recall that a scheme $S$ is essentially of finite type over a field $k$ if $S=\cup_{i=1}^{m}\operatorname{Spec}(A_{i})$, where $A_{i}$ are localizations of finite-type $k$-algebras.
An immediate consequence of Theorem 1.1 is the following:
Theorem 1.2.
Let $C\subset S$ be a hypersurface in a regular $2$-dimensional scheme $S$ which is essentially of finite type over a field $k$. Then the sequence of weighted blow-ups of Artin stacks introduced in Theorem 1.1, applied to $\mathcal{Z}=C$ and $\mathcal{W}=S$ and $T$ the trivial group, gives a stack-theoretic embedded resolution of singularities of $C\subset S$.
An application of Bergh’s destackification [B] provides a scheme-theoretic embedded resolution of $\mathcal{Z}\subset\mathcal{W}$ and of $[\mathcal{Z}/T]\subset[\mathcal{W}/T]$, see [AQS, Remark 5.6].
Rather than proving the result directly on higher dimensional schemes, we pass to the localization at the generic point of $\mathcal{Q}$ to reduce to the case of a singular hypersurface $C\subset S$ on a regular 2-dimensional scheme. This approach requires the following upgrade of [AQS, Theorem 1.1] to the case where $S$ is essentially of finite type over a field:
Theorem 1.3.
Let $C\subset S$ be a hypersurface in a regular $2$-dimensional scheme $S$ essentially of finite type over a field $k$. Let $q\in C$ be a singular closed point such that the reduction $C_{\rm red}$ of $C$ is singular at $q$.
  1. (1)
    There is a well-defined, unique center $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ of maximal invariant $(a_{1},a_{2})$ admissible for $C\subset S$ at $q$, where $a_{1}$ is the order of $C$ at $q$. Moreover, $J$ is stable under base change to separable field extensions of $k$.
  2. (2)
    Let $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$ be the associated reduced center. Its stack-theoretic weighted blow-up $S^{\prime}=Bl_{\bar{J}}(S)\to S$ is a proper birational morphism from a smooth Artin stack $S^{\prime}$ which is a global quotient of a variety by $\mathbb{G}_{m}$, in particular tame. (See [QR, Lemma 1.1.2 and Remark 3.2.10].)
  3. (3)
    The order of the proper transform of $C$ at every point of the blow-up lying over $q$ is strictly smaller than $a_{1}$.
Remark 1.4.
The assumption here that $S$ be essentially of finite type over a field is minimalistic — it is sufficient for our method to work. There is no doubt that one can further relax this assumption, but we opt to keep things simple in this paper.
Remark 1.5.
One could restrict attention to the case where the subscheme $C$ in Theorems 1.2, 1.3 and correspondingly $\mathcal{Z}$ in Theorem 1.1 is reduced. However the proof of Theorem 1.3 requires looking at the weighted normal cone of $C$, which is often non-reduced even when $C$ is. The generality we consider, with the reduction being singular, requires no additional effort.
This generality is also in line with model treatments of resolution of non-reduced schemes, such as [CJS, Definition 6.2]: a non-reduced scheme $X$ is quasi-regular if $X_{\operatorname{red}}$ is regular and $X$ normally flat along $X_{\operatorname{red}}$, namely the Hilbert–Samuel function is constant along $X_{\operatorname{red}}$ (see [CPSch]). A quasi-regular scheme is as resolved as a non-reduced scheme can possibly be.
But for a hypersurface, the Hilbert–Samuel function is automatically constant once $X_{\operatorname{red}}$ is regular. See also [CJS, Remarks 6.19, 6.20] and related discussion in [H1, Remark (1.13)]. Thus our singularity assumption is exactly that the subscheme $C$ or $\mathcal{Z}$ is not quasi-regular.
Acknowledgements
This research work was supported in part by funds from BSF grant 2022230, NSF grant DMS-2401358, and Simons Foundation Fellowship SFI-MPS-SFM-00006274. Dan Abramovich thanks the Hebrew University of Jerusalem for its hospitality during Spring 2025.
We thank André Belotto da Silva, Angélica Benito, Raymond van Bommel, Ana Bravo, Vincent Cossart, Santiago Encinas, Mike Montoro, James Myer, Michael Temkin and Jarosław Włodarczyk for discussions on the topic.
2. A brief introduction to Hironaka’s characteristic polyhedron
First, we give a short overview on Hironaka’s characteristic polyhedron [H1] in the special case of particular principal ideals which is the situation that we are interested in. Detailed introductions to the topic can be found in [CJS, Chapter 8 and appendix] as well as in [CSch], for instance.
Let $(\mathcal{O},\mathfrak{m},K)$ be a regular local ring with maximal ideal $\mathfrak{m}$ and residue field $K\simeq\mathcal{O}/\mathfrak{m}$. Let $f\in\mathfrak{m}$ be a non-zero element. Recall that the order of $f$ at $\mathfrak{m}$ is defined as
$\nu:=\nu_{\mathfrak{m}}(f):=\sup\{\alpha\in\mathbb{Z}_{+}\mid f\in\mathfrak{m}^{\alpha}\}.$
We make the assumption on $f$ that there exists a regular system of parameters $(x,y)=(x_{1},\ldots,x_{e},y)$ of $\mathcal{O}$ such that
(2.1) $f\equiv Y^{\nu}\mod\mathfrak{m}^{\nu+1}$
where $Y:=y\mod\mathfrak{m}^{2}$ denotes the image of $y$ in the graded ring $\operatorname{gr}_{\mathfrak{m}}(\mathcal{O})=\bigoplus_{a\geq 0}\mathfrak{m}^{a}/\mathfrak{m}^{a+1}$. In Remark 3.1 we address how we reduce to this situation for $1$-dimensional schemes essentially of finite type over a field. Since $\mathcal{O}$ is Noetherian, we have a finite expansion
$f=y^{\nu}+\sum_{(A,b)}c_{A,b}x^{A}y^{b}$
with $c_{A,b}\in\mathcal{O}^{\times}\cup\{0\}$ and multi-index notation $x^{A}=x_{1}^{A_{1}}\cdots x_{e}^{A_{e}}$, see [CP2, Proposition 2.1].
Definition 2.1.
  1. (1)
    The projected polyhedron of $f$ with respect to $(x;y)$ is defined as
    $\Delta(f;x;y)\ :=\ \operatorname{conv}\left(\left\{\frac{A}{\nu-b}+\mathbb{R}_{\geq 0}^{e}\ \Big{|}\ c_{A,b}\neq 0\right\}\right),$
    where $\operatorname{conv}(*)$ denotes the convex hull.
  2. (2)
    For a vertex $v\in\Delta(f;x;y)$ the initial form of $f$ at $v$ is given by
    $\operatorname{in}_{v}(f)\ :=\ Y^{\nu}+\sum_{(A,b):\frac{A}{\nu-b}=v}\overline{c_{A,b}}X^{A}Y^{b}\ \in\ K[X,Y],$
    where $\overline{c_{A,b}}$ are the images of the coefficients in the residue field $K\simeq\mathcal{O}/\mathfrak{m}$.
  3. (3)
    A vertex $v\in\Delta(f;x;y)$ is called solvable if $v\in\mathbb{Z}_{\geq 0}^{e}$ and there exists a (unique) $\lambda\in K\smallsetminus\{0\}$ such that $\operatorname{in}_{v}(f)=(Y-\lambda X^{v})^{\nu}$.
Illustrative pictures of the projected polyhedron and the initial form at a vertex can be found in [CJS, Examples 18.18, 18.23, and 18.31].
Suppose $v$ is a solvable vertex. Let $\epsilon\in\mathcal{O}^{\times}$ be any lift of $\lambda$ in $\mathcal{O}$, i.e., $\epsilon\equiv\lambda\mod\mathfrak{m}$. If we set $y^{\prime}:=y-\epsilon x^{v}$, then $v\notin\Delta(f;x;y^{\prime})$ and $\Delta(f;x;y)\supsetneq\Delta(f;x;y^{\prime})$ by [H1, Lemma (3.10)].
Example 2.2.
Let $\mathcal{O}=K[[x,y]]$ be of dimension two for a field $K$ of characteristic 2 and consider the polynomial $f=y^{2}+\lambda x^{4}+x^{5}$ with $\lambda\in K$. The vertex of the projected polyhedron $\Delta(f;x;y)$ is $v=2$.
If $\lambda\in K^{2}$ is a square, then $v$ is solvable via the change $y^{\prime}:=y+\mu x^{2}$ with $\mu\in K$ such that $\mu^{2}=\lambda$. The resulting polyhedron $\Delta(f;x;y^{\prime})$ has vertex $v^{\prime}=5/2$ which is not solvable.
On the other hand, if $K$ is imperfect and if $[K^{2}(\lambda):K^{2}]=2$, i.e., $\lambda$ is not a square, then $v=2$ is not solvable.
For a non-empty convex subset $\Delta\subseteq\mathbb{R}_{\geq 0}^{e}$, we define the number
$\delta(\Delta):=\inf\{v_{1}+\cdots+v_{e}\mid(v_{1},\ldots,v_{e})\in\Delta\}.$
Furthermore we set $\delta(\emptyset):=\infty$ and abbreviate $\delta(f;x;y):=\delta(\Delta(f;x;y))$.
Let us point out that $\delta(f;x;y)=\infty$ if and only if $f=\epsilon y^{\nu}$ for some unit $\epsilon\in\mathcal{O}$. In this case the reduction of $\{f=0\}$ is regular.
Theorem 2.3.
Let $f\in\mathfrak{m}$ be an element such that (2.1) holds.
  1. (1)
    There is an element $z\in\mathcal{O}$ such that $(x,z)$ is a regular system of parameters for $\mathcal{O}$ and $\Delta(f;x;z)$ is minimal with respect to inclusion among all possible $\Delta(f;x;y)$ with $(x,y)$ a regular system of parameters for $\mathcal{O}$.
  2. (2)
    The number $\delta(f;x;z)$ is an invariant of the singularity $\mathcal{O}/(f)$.
The minimal polyhedron $\Delta(f;x):=\Delta(f;x;z)$ appearing in Theorem 2.3(1) is called the characteristic polyhedron of $f$. We emphasize that this is an ad-hoc definition. Originally, $\Delta(f;x)$ is defined as the intersection over all possible projected polyhedra $\Delta(f;x;y)$ with varying $y$.
Proof of Theorem 2.3.
This follows from a combination of results from [H1, CP1, CJSch]: Hironaka proved that in the $\mathfrak{m}$-adic completion $\widehat{\mathcal{O}}$ there is an element $\widehat{z}$ with $\Delta(f;x;\widehat{z})=\Delta(f;x)$ in [H1, Theorems (3.17) and (4.8)]. In [CP1] the existence of $z\in\mathcal{O}$, not necessitating completion, was shown for the given setting. Generalizations followed in [CSch]. The second part on the invariant is proven in [CJSch, Corollary B(3)]. ∎
Definition 2.4.
Let $f\in\mathcal{O}$ be as above and let $z$ be as in Theorem 2.3. Set $\delta:=\delta(f;x;z)$. The $\delta$-initial form of $f=\sum_{(A,b)}c_{A,b}x^{A}z^{b}$ is defined as the initial form of $f$ at the vertex $\delta$, which can be rewritten as
$\operatorname{in}_{\delta}(f)\ =\ \sum_{(A,b)\,:\,\frac{|A|}{\delta\nu}+\frac{b}{\nu}=1}\overline{c_{A,b}}X^{A}Z^{b}.$
We introduce the following lift to $\mathcal{O}$ of the $\delta$-initial form of $f$:
$F_{\delta}\ :=\ \sum_{(A,b)\,:\,\frac{|A|}{\delta\nu}+\frac{b}{\nu}=1}c_{A,b}x^{A}z^{b}\in\mathcal{O}.$
Observe that the definition of the lift depends on the choice of lifts for the coefficients of the initial forms from $K=\mathcal{O}/\mathfrak{m}$ to $\mathcal{O}$. Since they originate from certain elements $c_{A,b}\in\mathcal{O}^{\times}\cup\{0\}$, we choose them as the lifts.
Example 2.5.
  1. (1)
    Let $k$ be an imperfect field of characteristic $p$ and $a\in k\smallsetminus k^{p}$. Consider
    $f\,=\,y^{p}-x_{1}^{m}(x_{1}^{p}+a)^{pm+1}-ax_{2}^{pm}\ \in\ k[x_{1},x_{2},y]$
    with $m>p$ coprime to $p$.
    At the closed point $q$ given by the maximal ideal $(x_{1},x_{2},y)$ the residue field is equal to $k$. The vertices of $\Delta(f;x;y)$ are $v_{1}:=(\frac{m}{p},0)$ and $v_{2}:=(0,m)$. Observe that $\operatorname{in}_{v_{2}}(f)=Y^{p}-aX_{2}^{pm}$ and since $a\notin k^{p}$, the vertex $v_{2}$ is not solvable. Thus, none of the vertices is solvable and $\Delta(f;x;y)=\Delta(f;x)$ is the characteristic polyhedron (at $q$).
    Let us look at the point $\widetilde{q}$ corresponding to the maximal ideal $(x_{1}^{p}+a,x_{2},y)$. The residue field at this point is $K:=k[x_{1}]_{(x_{1}^{p}+a)}/(x_{1}^{p}+a)$. Possible local parameters at $\widetilde{q}$ are $(u_{1}:=x_{1}^{p}+a,x_{2},y)$. By substituting $a=u_{1}-x_{1}^{p}$, $f$ can be rewritten as
    $f\,=\,y^{p}-x_{1}^{m}u_{1}^{pm+1}-u_{1}x_{2}^{pm}+x_{1}^{p}x_{2}^{pm}.$
    Taking into account that $x_{1}$ is invertible at $\widetilde{q}$, the vertices of the corresponding projected polyhedron $\Delta(f;u_{1},x_{2};y)$ are $\widetilde{v_{1}}:=(m+\frac{1}{p},0)$ and $\widetilde{v_{2}}:=(0,m)$. This time $\widetilde{v_{2}}$ is solvable since $\operatorname{in}_{\widetilde{v_{2}}}(f)=Y^{p}+\overline{x_{1}}^{p}X_{2}^{pm}=(Y+\overline{x_{1}}X_{2}^{m})^{p}$ where $\overline{x_{1}}$ denotes the image of $x_{1}$ in the residue field $K$. Introducing $z:=y+x_{1}x_{2}^{m}$ the vertices of $\Delta(f;u_{1},x_{2};z)$ are $\widetilde{v_{1}}$ and $\widetilde{v_{3}}:=(\frac{1}{p},m)$. Both vertices are not solvable and hence $\Delta(f;u_{1},x_{2};z)=\Delta(f;u_{1},x_{2})$ is the characteristic polyhedron (at $\widetilde{q}$).
  2. (2)
    Let $k$ be a perfect field of positive characteristic $p$ and let $\mathcal{O}$ be the localization of $k[x,y,z]$ at the ideal $(x,y)$. Let $m\in\mathbb{Z}_{+}$ with $m>p$ and
    $f\,=\,y^{2p}+x^{m}y^{p}z^{m+p}+x^{2m}z^{p+1}+x^{2m-1}y\ \in\ \mathcal{O}.$
    Notice that $(x,y)$ is a regular system of parameters for $\mathcal{O}$ and $z$ is invertible in $\mathcal{O}$. The projected polyhedron $\Delta(f;x;y)=\Delta(f;x)$ coincides with the characteristic polyhedron since the vertex is $\delta=\frac{m}{p}$ which is not solvable.
    We have $\operatorname{in}_{\delta}(f)=Y^{2p}+X^{m}Y^{p}\overline{z}^{m+p}+X^{2m}\overline{z}^{p+1}$, where $\overline{z}$ denotes the image of $z$ in the residue field $\mathcal{O}/(x,y)$.
3. Invariant, parameters, and center
In the next step, we define the invariant in the setting of Theorem 1.3 and deduce from this the designated center for the weighted blow-up.
Let $C\subset S$ be a hypersurface in a regular $2$-dimensional scheme $S$ which is essentially of finite type over a field $k$ and let $q\in C$ be a singular closed point such that $C_{\rm red}$ is singular at $q$. Let $(\mathcal{O},\mathfrak{m},K)$ be the local ring of $C$ at $q$ and let $f\in\mathfrak{m}$ be an element locally at $q$ defining $C$. Sometimes, we will abuse notation by writing $C=\{f=0\}$.
Remark 3.1.
Let us address the situation if (2.1) does not hold. Then $\operatorname{in}_{\mathfrak{m}}(f):=f\mod\mathfrak{m}^{\nu+1}\in\operatorname{gr}_{\mathfrak{m}}(\mathcal{O})$ cannot be expressed as polynomial in a single variable. For readers familiar with Hironaka’s notion of the directrix (for example, see [CJS, Definitions 2.8 and 2.26]), the latter means that the dimension of the directrix is zero. Due to a result of Hironaka [H2, Theorem 2] which was generalized by Mizutani [M] (see also [CJS, Theorem 3.14]) blowing up the closed point $q$ will lead in this situation to a strict decrease of the order at points lying above $q$.
In the remainder of the article we include the discussed case by setting $\delta:=1$ and $\operatorname{in}_{\delta}(f):=\operatorname{in}_{\mathfrak{m}}(f)$ if (2.1) does not hold.
Definition 3.2.
Let $q\in C\subset S$ be as before, locally given as $q\in\{f=0\}\subset\operatorname{Spec}(\mathcal{O})$. The invariant of $C$ at $q$ is defined as
$\operatorname{inv}_{C}(q)\,:=\,\operatorname{inv}_{f}(q)\,:=\,(a_{1},a_{2})\,:=\,(\nu,\delta\nu)\ \,\in\ \,\Gamma\,:=\,\mathbb{Z}_{\geq 0}\times\frac{1}{a_{1}!}\mathbb{Z}_{\geq 0},$
where $\nu$ is the order of $C$ at $q$, $\delta\in\mathbb{Q}_{+}$ is the number introduced in Theorem 2.3(2) (resp. Remark 3.1), and $\Gamma$ is equipped with the lexicographical ordering.
Recall that our assumption, $C_{\rm red}$ is singular at $q$, implies that $\delta:=\delta(f;x;z)<\infty$, see the paragraph just before Theorem 2.3.
Using parameters $(z,x)$ with the property $\Delta(f;x;z)=\Delta(f;x)$ we determine the center of the weighted blow-up in terms of monomial valuations. While this definition depends on choices, we show in Theorem 3.5 that the center is unique. Despite working in a more general setup, the proof follows the same lines as [AQS, Theorem 4.2].
Definition 3.3.
Let $(\mathcal{O},\mathfrak{m},K)$ be a regular local ring of dimension $2$ which is essentially of finite type over a field $k$.
  1. (1)
    For any parameters $(x_{1},x_{2})$ of $\mathcal{O}$ and any positive rational numbers $a_{1},a_{2}\in\mathbb{Q}_{+}$, we define $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ to be the monomial valuation $v_{J}\colon\mathcal{O}\smallsetminus\{0\}\to\mathbb{Q}$ given by
    $v_{J}(x_{1})=\frac{1}{a_{1}},\quad v_{J}(x_{2})=\frac{1}{a_{2}}.$
  2. (2)
    For a non-zero element $f\in\mathcal{O}\smallsetminus\{0\}$, we say $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ is admissible for $f$ if we have $v_{J}(f)\geq 1$.
Remark 3.4.
  1. (1)
    Since $v_{J}$ is a monomial valuation, we have: For any $g\in\mathcal{O}\smallsetminus\{0\}$ with expansion $g(x_{1},x_{2})=\sum\limits_{(i_{1},i_{2})}d_{i_{1},i_{2}}x_{1}^{i_{1}}x_{2}^{i_{2}}$ for $d_{i_{1},i_{2}}\in\mathcal{O}^{\times}\cup\{0\}$, the following holds:
    $v_{J}(g)=\min\left\{\frac{i_{1}}{a_{1}}+\frac{i_{2}}{a_{2}}\ \Big{|}\ d_{i_{1},i_{2}}\neq 0\right\}.$
    In particular, we have $v_{J}(x_{1}^{a_{1}})=v_{J}(x_{2}^{a_{2}})=1$.
  2. (2)
    Observe that $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ is admissible for a given $f$ if and only if $f\in(x_{1}^{a_{1}},x_{2}^{a_{2}})^{{\operatorname{Int}}}$.
  3. (3)
    For resolution of $1$-dimensional schemes the local perspective via valuations is enough. In general, for instance in Theorem 1.1, one has to describe centers globally, as for example in [ATW, Section 2.2] via valuative $\mathbb{Q}$-ideals.
Theorem 3.5.
Let $q\in C=\{f=0\}\subset\operatorname{Spec}(\mathcal{O})$ be as described at the beginning of this section. Let $(a_{1},a_{2}):=(\nu,\nu\delta)$ and let $(x_{1},x_{2})$ be a regular system of parameters for $\mathcal{O}$ such that $\Delta(f;x_{2};x_{1})=\Delta(f;x_{2})=[\delta;\infty[$. Then $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ is the unique admissible center for $f$ for which $(a_{1},a_{2})$ with $a_{1}\leq a_{2}$ attains its maximum with respect to the lexicographical order. Moreover, $J$ is stable under base change to separable field extensions of $k$.
We note that this theorem implies Theorem 1.3(1).
Proof.
Since the proof is analogous to the one for [AQS, Theorem 4.2], we outline the steps and refer for further details to [AQS]. Our goal is to show
(3.1) $\max_{\geq_{\rm lex}}\{(b_{1},b_{2})\in\mathbb{Q}_{+}^{2}\mid\exists\,(y_{1},y_{2}):(y_{1}^{b_{1}},y_{2}^{b_{2}})\mbox{ admissible center for }f\}=(a_{1},a_{2}).$
First, the left hand side of Equation (3.1) is $\geq(a_{1},a_{2})$ since $J$ is admissible by the definition of to the characteristic polyhedron.
Secondly, $a_{1}$ is unique: If $(y_{1}^{b_{1}},y_{2}^{b_{2}})$ is another admissible center with $b_{1}\leq b_{2}$, we must have $b_{1}\leq a_{1}$ since $a_{1}=\nu_{\mathfrak{m}}(f)$ which implies the existence of a monomial $y_{1}^{a}y_{2}^{\nu-a}$ with coefficient a unit in the expansion of $f$.
Thirdly, $a_{2}$ is unique: Let $(y_{1}^{\nu},y_{2}^{b_{2}})$ be an admissible center. Using the definition of the characteristic polyhedron and $\delta(f;y_{2};y_{1})\leq\delta$, it follows that we have to have $b_{2}\leq\nu\delta=a_{2}$ which implies Equation (3.1).
Finally, we show that the center is unique: Let $J_{x}=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ and $J_{y}=(y_{1}^{a_{1}},y_{2}^{a_{2}})$ be two admissible centers for $f$. If $a_{1}=a_{2}$ then $v_{J_{x}}=v_{J_{y}}=a_{1}\nu_{m}$ and the desired equality $J_{x}=J_{y}$ holds. Hence, suppose $a_{1}<a_{2}$.
  1. (a)
    $x_{2}=y_{2}$ : By construction, we may vary $x_{2}$ for the given $x_{1}$ as long as we have $\mathfrak{m}=(x_{1},\widetilde{x_{2}})$ for the different choice. The analogous statement is true for $y_{2}$. Since $\mathbb{P}(\mathfrak{m}/\mathfrak{m}^{2})$ has at least 3 points, we may assume without loss of generality that $x_{2}=y_{2}$. Further, this provides $v_{J_{y}}(x_{2})=v_{J_{x}}(x_{2})=\frac{1}{a_{2}}$.
  2. (b)
    $v_{J_{y}}(x_{1})=v_{J_{x}}(y_{1})=\frac{1}{a_{1}}$ : If $x_{1}$ and $y_{1}$ differ by a unit factor, the assertion clearly holds. Therefore, consider an expansion of $x_{1}$ in the regular system of parameter $(y_{1},x_{2})$, say $x_{1}=\epsilon y_{1}+\mu x_{2}^{A}$ for $A\in\mathbb{Z}_{+},\epsilon,\mu\in\mathcal{O}^{\times}$. Since $v_{J_{y}}$ is a monomial valuation, this provides $v_{J_{y}}(x_{1})\leq v_{J_{y}}(y_{1})=\frac{1}{a_{1}}$.
    Assume the inequality is strict, $v_{J_{y}}(x_{1})<\frac{1}{a_{1}}$. Expand $f=\sum_{(i_{1},i_{2})}d_{i_{1},i_{2}}x_{1}^{i_{1}}x_{2}^{i_{2}}$. Since $J_{x}$ is admissible for $f$, we have $\frac{i_{2}}{a_{2}}\geq\frac{a_{1}-i_{1}}{a_{1}}$ for every $(i_{2},i_{2})$ such that $d_{i_{1},i_{2}}\in\mathcal{O}^{\times}$ is a unit. Moreover, we have an equality if and only if $i_{1}=a_{1},i_{2}=0$. Using the non-Archimedean property, it follows that $v_{J_{y}}(f)=a_{1}v_{J_{y}}(x_{1})<1$. This contradicts the assumption that $J_{y}$ is an admissible center for $f$ and therefore we obtain $v_{J_{y}}(x_{1})=\frac{1}{a_{1}}$. Applying the analogous arguments leads to $v_{J_{x}}(y_{1})=\frac{1}{a_{1}}$.
  3. (c)
    $J_{x}=J_{y}$ : This follows since $(x_{1}^{ma_{1}},x_{2}^{ma_{2}})^{\operatorname{Int}}=(y_{1}^{ma_{1}},y_{2}^{ma_{2}})^{\operatorname{Int}}$ for every $m\in\mathbb{Z}_{\geq 0}$.
It remains to prove that $J$ is stable under base change to separable field extensions of $k$. Let $k^{\prime}/k$ be a separable extension. Using [CP2, Prop. 2.15] (applied for $h:=\operatorname{in}_{\delta}(f)$ considered as element in $S[X_{1}]$ for $S:=K[X_{2}]_{(X_{2})}$ and $\widetilde{S}:=S\otimes_{k}k^{\prime}$) it can be deduced that the characteristic polyhedron does not change under the base change $k^{\prime}/k$ and thus $J$ remains stable.
Since our given situation is simpler, we provide a direct argument: We have to study the behavior of $\phi:=\operatorname{in}_{\delta}(f)\in K[X_{1},X_{2}]$ under separable extensions. Let $\phi_{s}$ be the element $\phi$ considered as element in $(K\otimes_{k}k^{\prime})[X_{1},X_{2}]$. We have to show that $\phi_{s}$ is not a $\nu$-th power, $\phi_{s}\neq(X_{1}+\alpha X_{2}^{\delta})^{\nu}$ for $\alpha\in(K\otimes_{k}k^{\prime})\smallsetminus\{0\}$. If $\delta\notin\mathbb{Z}_{+}$, this is certainly not possible. Hence let $\delta\in\mathbb{Z}_{+}$ and assume $\phi_{s}=(X_{1}+\alpha X_{2}^{\delta})^{\nu}$. Write $\nu=p^{m}\nu^{\prime}$ for $\nu^{\prime}$ prime to $p$ if $\operatorname{char}(k)=p>0$ and $\nu=\nu^{\prime},p^{m}=1$ else. Since the extension is separable, we have to have $\phi=\psi^{p^{m}}$ for some $\psi\in K[X_{1},X_{2}]$. Clearly, $\nu^{\prime}\geq 2$ in this case and $\psi_{s}=(X_{1}+\alpha X_{2}^{\delta})^{\nu^{\prime}}$. We have $\psi=X_{1}^{\nu^{\prime}}+\sum_{i=1}^{\nu^{\prime}}c_{i}X_{1}^{\nu^{\prime}-i}X_{2}^{\delta i}$ for $c_{i}\in K$. Let $Y_{1}:=X_{1}+\frac{c_{1}}{\nu^{\prime}}X_{2}^{\delta}$. Substituting $X_{1}=Y_{1}-\frac{c_{1}}{\nu^{\prime}}X_{2}^{\delta}$ provides $\psi=Y_{1}^{\nu^{\prime}}+\sum_{i=2}^{\nu^{\prime}}c_{i}^{\prime}Y_{1}^{\nu^{\prime}-i}X_{2}^{\delta i}\neq Y_{1}^{\nu^{\prime}}$ for certain $c_{i}^{\prime}\in K$. Observe that $J$ remains stable under this substitution and that $c_{1}^{\prime}=0$. The last condition is stable by the base change with $k^{\prime}$, but this contradicts $\psi_{s}=(X_{1}+\alpha X_{2}^{\delta})^{\nu^{\prime}}$. ∎
In order to draw a connection to the weighted normal cone, we need to compare $f$ and $F_{\delta}$ (Definition 2.4). Note that $F_{\delta}$ is the sum over all monomials of valuation $1$ in an expansion of $f$. To emphasize the relation to the valuation $v_{J}$, we introduce the notation $f_{J,1}:=F_{\delta}.$ Hence, we have
(3.2) $f=f_{J,1}+\widetilde{h},\quad\mbox{ where }\widetilde{h}\in\mathcal{O}\mbox{ with }v_{J}(\widetilde{h})>1.$
Corollary 3.6.
The reduction of $\{f_{J,1}=0\}$ is singular at $q$ and $\operatorname{inv}_{f_{J,1}}(q)=(a_{1},a_{2})$ with the same center $(x_{1}^{a_{1}},x_{2}^{a_{2}})$ as for $f$.
Proof.
First, we have $\operatorname{in}_{\delta}(f_{J,1})=\operatorname{in}_{\delta}(f)$. This implies that the vertex $\delta\in\Delta(f_{J,1};x_{2};x_{1})$ is not solvable. In particular, $f_{J,1}$ is not equal to the power of a regular parameter which implies the claim on the reduction of $\{f_{J,1}=0\}$.
Furthermore, the characteristic polyhedra of $f_{J,1}$ and $f$ coincide, $\Delta(f_{J,1};x_{2};x_{1})=\Delta(f_{J,1};x_{2})=[\delta;\infty[=\Delta(f;x_{2})$. So, the designated center for $f_{J,1}$ is $(x_{1}^{a_{1}},x_{2}^{a_{2}})$ as well. ∎
4. Weighted blow-ups and proper transform
Now that we have a candidate for the center, let us recall the concept of a stack-theoretic weighted blow-up. A general reference for them is [QR]. Weighted blow-ups are very closely connected to Włodarczyk’s notion of cobordant blow-ups [W], where a torus action is taken into account instead of taking the stack-theoretic quotient.
Definition 4.1 ([ATW, Section 2.4]).
For $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$, the associated reduced center is $\bar{J}:=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$, with $w_{1},w_{2}\in\mathbb{Z}_{+}$ relatively prime integers such that
$\ell\cdot\left(\frac{1}{w_{1}},\frac{1}{w_{2}}\right)\,=\,(a_{1},a_{2})$
for $\ell:=a_{1}w_{1}=a_{2}w_{2}$ the least multiple of $a_{1},a_{2}$ in $\mathbb{Z}_{+}$.
We introduce the weighted blow-up in terms of extended Rees algebras [R]. For $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ with reduced center $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$ and $n\in\mathbb{Z}$ set $\mathcal{I}_{n}:=\{g\in\mathcal{O}\mid v_{\bar{J}}(g)\geq n\}.$ Notice that that $\mathcal{I}_{n}=\mathcal{O}$ for $n\leq 0$. The extended Rees algebra associated to $\bar{J}$ is the $\mathcal{O}$-algebra
$\tilde{R}\,=\,\bigoplus_{n\in\mathbb{Z}}\mathcal{I}_{n}.$
This local construction becomes global via $\mathcal{I}_{n}=\mathcal{O}_{S}$ away from $q$.
By [QR, Proposition 5.2.2], $\tilde{R}$ has the following local presentation:
(4.1) $\tilde{R}\,=\,\mathcal{O}[s,x_{1}^{\prime},x_{2}^{\prime}]/(x_{1}-s^{w_{1}}x_{1}^{\prime},x_{2}-s^{w_{2}}x_{2}^{\prime}),$
for $s\in\mathcal{I}_{-1}$ the element $1\in\mathcal{O}$, and $x_{i}^{\prime}\in\mathcal{I}_{w_{i}}$ corresponding to $x_{i}$ in degree $w_{i}$, $i\in\{1,2\}$. Here, $s$ is the exceptional variable, defining the exceptional divisor once the blow-up is defined. Moreover, $x_{i}^{\prime}$ is called the transformed variable corresponding to $x_{i}$, for $i\in\{1,2\}$.
Construction 4.2 (Degenerating to the weighted normal cone, see [W, QR]).
Let $B=\operatorname{Spec}_{S}(\tilde{R})$ be the spectrum relative to the scheme $S$ of the quasi-coherent sheaf of $\mathcal{O}_{S}$-algebras $\tilde{R}$. Besides from its structure morphism $B\to S$, $B$ is also equipped with a morphism $\varphi\colon B\to\mathbb{A}_{k}^{1}=\operatorname{Spec}k[s]$. Using the latter $B$ becomes the degeneration of $S$ to the weighted normal cone of $\bar{J}$. Using Equation (4.1), we have:
  • The fiber of $\varphi$ over $s=1$ is isomorphic to $S$ since $x_{1}=x_{1}^{\prime},x_{2}=x_{2}^{\prime}$.
  • The fiber of $\varphi$ over $s=0$ lies over the point $(x_{1},x_{2})=(0,0)$ and it is isomorphic to $\operatorname{Spec}k[x_{1}^{\prime},x_{2}^{\prime}]$ with the natural action of $\mathbb{G}_{m}$ with weights $w_{i}$ on $x_{i}^{\prime}$, making it the weighted normal cone at $(0,0)\in S$.
Observe that we have the following action of $\mathbb{G}_{m}=\operatorname{Spec}k[t,t^{-1}]$ on $B$
(4.2) $t\cdot(s,x_{1}^{\prime},x_{2}^{\prime})\quad=\quad(t^{-1}s,t^{w_{1}}x_{1}^{\prime},t^{w_{2}}x_{2}^{\prime})$
stabilizing $V(\oplus_{n>0}\mathcal{I}_{n})=V(x_{1}^{\prime},x_{2}^{\prime})$.
Definition 4.3 (Weighted blow-up).
Let $B=\operatorname{Spec}_{S}(\tilde{R})$ be as in Construction 4.2. The weighted blow-up of $S$ along $\bar{J}$ is defined as the Artin stack
$\pi\colon\operatorname{Bl}_{\bar{J}}(S):=[B_{+}/\mathbb{G}_{m}]\to S,$
where $B_{+}:=B\smallsetminus V(x_{1}^{\prime},x_{2}^{\prime})$, with natural $\mathbb{G}_{m}$-action determined by (4.2) and the morphism $\pi$ is induced by the structure morphism $B\to S$.
Observe that $\operatorname{Bl}_{\bar{J}}(S)$ is covered by two charts, namely where $x_{1}^{\prime}$ (resp. $x_{2}^{\prime}$) is invertible in which case we also write $x_{1}^{\prime}\neq 0$ (resp. $x_{2}^{\prime}\neq 0$).
By [QR, Lemma 1.1.2 and Remark 3.2.10] the weighted blow-up fulfills the following properties. Note that this proves Theorem 1.3(2).
Lemma 4.4.
Let $\pi\colon\operatorname{Bl}_{\bar{J}}(S)=[B_{+}/\mathbb{G}_{m}]\to S$ be the weighted blow-up of $S$ along the reduced center $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$ coming from $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$. Then:
  1. (1)
    All stabilizers of the Artin stack $\operatorname{Bl}_{\bar{J}}(S)$ are naturally subgroups of $\mathbb{G}_{m}$. In particular, $\operatorname{Bl}_{\bar{J}}(S)$ is a tame Artin stack.
  2. (2)
    The dimension of the quotient $\operatorname{Bl}_{\bar{J}}(S)$ is two and the morphism $\pi$ is an isomorphism outside the exceptional divisor, which is defined on the smooth covering of $B_{+}$ by $s=0$.
In particular, the weighted blow-up $\pi$ is birational.
  1. (3)
    At the point $s=x_{1}^{\prime}=0$ the stabilizer group is $\boldsymbol{\mu}_{w_{2}}$; while at the point $s=x_{2}^{\prime}=0$ the stabilizer group is $\boldsymbol{\mu}_{w_{1}}$. Here, $\boldsymbol{\mu}_{m}$ denotes the group of $m$-th roots of unity for $m\in\mathbb{Z}_{+}$.
Let us next have a look at how an element $f\in\mathcal{O}$ transforms along a weighted blow-up. Recall that we have given a singular point $q$ on a hypersurface $C=\{f=0\}$ in a regular $2$-dimensional scheme which is essentially of finite type over a field $k$.
Consider an expansion of $f$ as in Equation (3.2), $f=f_{J,1}(x_{1},x_{2})+\widetilde{h}(x_{1},x_{2})$. Recall that $v_{\bar{J}}(f)=v_{\bar{J}}(f_{J,1})=\ell<v_{\bar{J}}(\widetilde{h})$ where $\ell\in\mathbb{Z}_{+}$ is the integer introduced in Definition 4.1. Using the presentation of $\tilde{R}$ given in Equation (4.1), the pullback of $f$ to $B$ is
$f(x_{1},x_{2})\ =\ f(s^{w_{1}}x_{1}^{\prime},s^{w_{2}}x_{2}^{\prime})\ =\ s^{\ell}f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})+s^{\ell+1}\widetilde{h}^{\prime}(s,x_{1}^{\prime},x_{2}^{\prime})$
where $\widetilde{h}^{\prime}(s,x_{1}^{\prime},x_{2}^{\prime})$ is defined through $\widetilde{h}(s^{w_{1}}x_{1}^{\prime},s^{w_{2}}x_{2}^{\prime}):=s^{v_{\bar{J}}(\widetilde{h})-(\ell+1)}\widetilde{h}^{\prime}(s,x_{1}^{\prime},x_{2}^{\prime})$.
Definition 4.5.
The proper transform of $f$ on $B$ is defined as the element $f^{\prime}$ that we obtain after factoring $s^{\ell}$ from the pullback of $f$ to $B$,
$f^{\prime}(s,x_{1}^{\prime},x_{2}^{\prime})\ :=\ s^{-\ell}f(s^{w_{1}}x_{1}^{\prime},s^{w_{2}}x_{2}^{\prime})\ =\ f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})+s\,\widetilde{h}^{\prime}(s,x_{1}^{\prime},x_{2}^{\prime}).$
Example 4.6.
Consider again Example 2.5(2), and assume for simplicity that $p\nmid m$. We have determined $\nu=2p$ and $\delta=m/p$, giving $(a_{1},a_{2})=(2p,2m)$ and, keeping $(x_{1},x_{2})=(y,x)$ we have $J=\left(y^{2p},x^{2m}\right)$. We now write $f=f_{J,1}(y,x)+\widetilde{h}(y,x)$ with
$f_{J,1}(y,x)\ =\ y^{2p}+x^{m}y^{p}z^{m+p}+x^{2m}z^{p+1}\qquad\text{and}\qquad\widetilde{h}(y,x)\ =\ x^{2m-1}y.$
Since $p\nmid m$ we have $\ell=2mp$. We rescale to obtain $\bar{J}=\left(y^{1/m},x^{1/p}\right)$, so that $x=s^{p}x^{\prime},y=s^{m}y^{\prime},$ giving
$f^{\prime}(y^{\prime},x^{\prime})\quad=\quad y^{\prime 2p}+x^{\prime m}y^{\prime p}z^{m+p}+x^{\prime 2m}z^{p+1}\quad+\quad s^{m-p}x^{\prime 2m-1}y^{\prime}.$
The argument in the Section 5 below guarantees that its order along the exceptional locus $s=0$ is $<2p$. This follows from the fact that the term $y^{\prime 2p}+x^{\prime m}y^{\prime p}z^{m+p}+x^{\prime 2m}z^{p+1}$ is not of the form $g^{2p}$ and that $x^{\prime}$ or $y^{\prime}$ is invertible (and also remember that $z$ is invertible).
5. Proofs
Our final section is devoted to proving Theorem 1.3(3) and Theorem 1.1. Applying the same arguments as explained in [AQS, after Defintion 6.1], Corollary 3.6 can be used to deduce that the invariant on any fiber of $\varphi|_{B_{+}}\colon B_{+}\to\mathbb{A}_{k}^{1}$ (see Construction 4.2) drops strictly. As in [AQS], we show the strong statement that the order decreases strictly after the designated weighted blow-up.
The order drops: Proof of Theorem 1.3(3).
Since the order is functorial for smooth morphisms, showing that the order drops on $B_{+}$ implies that it drops on the quotient ${\operatorname{Bl}}_{\bar{J}}(S)=[B_{+}/\mathbb{G}_{m}]$. Note that $\nu_{\mathfrak{m}}(f)\leq\nu^{\log}_{\mathfrak{m}}(f)$, where the logarithmic order $\nu^{\log}_{\mathfrak{m}}(f)$ is defined with respect to the exceptional divisor $\{s=0\}$: it is the order where $s\neq 0$ and otherwise it is the order of the restriction to $\{s=0\}$, namely the order of $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})$.
When $s\neq 0$ we simply are computing the order of $f$ away from $q=\{x_{1}=x_{2}=0\}$. We note that near $q$ this order is $<\nu=a_{1}$ since $f$ is not a pure $\nu$-th power.
When $s=0$ we are computing the order of $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})$ away from $\{x_{1}^{\prime}=x_{2}^{\prime}=0\}$. But once again, $f_{J,1}(x_{1}^{\prime},x_{2}^{\prime})=x_{1}^{\prime\nu}+\cdots$ is not a pure $\nu$-th power, see Corollary 3.6. Thus its order of vanishing at any point on this locus is $<\nu=a_{1}$, as needed. ∎
Using this result, we can now prove its equivariant generalization.
Proof of Theorem 1.1.
Consider the localizations of $\mathcal{Z}\subset\mathcal{W}$ at the generic point $q:=\eta_{\mathcal{Q}}$ of $\mathcal{Q}$, namely $C:=\operatorname{Spec}(\mathcal{O}_{\mathcal{Z},q})$ and $S:=\operatorname{Spec}(\mathcal{O}_{\mathcal{W},q})$. Then $S$ is a 2-dimensional regular scheme that is essentially of finite type over $k$, and $C$ is cut out by a local equation $\{f=0\}$ in $S$. Note that the residue field of the closed point $q\in S$ is the function field $K(\mathcal{Q})$ and hence it is not necessarily perfect even if $k$ was assumed to be perfect.
We apply part (1) of Theorem 1.3 to $C\subset S$ at $q$. There is a unique center $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ of maximal invariant $(a_{1},a_{2})$ that is admissible for $C\subset S$ at $q$, where $a_{2}=a_{1}\delta$, with $\delta$ the vertex of the characteristic polyhedron of $f$. Let $q\in U\subset\mathcal{Q}$ be the largest open subset on which $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ extends. In what follows, we show $U=\mathcal{Q}$:
  1. (i)
    Firstly, letting $k^{\operatorname{sep}}$ denote the separable closure of $k$, we claim $U$ is stable under $T(k^{\operatorname{sep}})$. Indeed, $J$ remains the unique center of maximal invariant $(a_{1},a_{2})$ that is admissible for $C\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})\subset S\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})$ at $q$. On the other hand, for every $t\in T(k^{\operatorname{sep}})$, $t\cdot J$ is also a center on $S\times_{\operatorname{Spec}(k)}\operatorname{Spec}(k^{{\operatorname{sep}}})$ that satisfies the same properties, so by the uniqueness of $J$, $t\cdot J=J$. Thus, $J=t\cdot J$ extends to $t\cdot U\subset\mathcal{Q}$, so $t\cdot U\subset U$.
  2. (ii)
    Secondly, we claim that in fact $U$ is stable under $T$. To do so, we instead show that $F=\mathcal{Q}\smallsetminus U$ is stable under $T$, or equivalently, that the preimage of $F$ under the $T$-action $\alpha\colon T\times\mathcal{Q}\to\mathcal{Q}$ contains $T\times F$. By (i), $\alpha^{-1}(F)\supset\{t\}\times F$ for every $t\in T(k^{\operatorname{sep}})$. Since $T$ is integral and geometrically reduced, the points in $T(k^{\operatorname{sep}})$ are dense in $T$. Thus, after taking closures, we deduce $\alpha^{-1}(F)\supset T\times F$ as desired.
Thus, $U\subset\mathcal{Q}$ are both non-empty $T$-orbits, so $U=\mathcal{Q}$. Hence, $J$ extends to $\mathcal{Q}$, and moreover by (i), that extension to $\mathcal{Q}$ is $T$-invariant.
Next, we claim that $J=(x_{1}^{a_{1}},x_{2}^{a_{2}})$ is the unique $T$-invariant center of maximal invariant that is admissible for $\mathcal{Z}\subset\mathcal{W}$ along $\mathcal{Q}$. Note that a $T$-invariant center is necessarily defined by at most 2 local parameters at all points. To see the claim, any other such center $J^{\prime}$ necessarily of invariant $(a_{1},a_{2})$ would localize to a center of maximal invariant that is admissible for $C\subset S$ at $q$. By the uniqueness of $J$ in Theorem 1.3, $J$ agrees with $J^{\prime}$ at $q$, and hence on a non-empty open subset of $\mathcal{Q}$. Since $\mathcal{Q}$ is an orbit under the $T$-action on $\mathcal{W}$, and both $J$ and $J^{\prime}$ are $T$-invariant, we deduce that $J=J^{\prime}$. This completes the proof of part (1) of Theorem 1.1.
We now prove the remaining parts of Theorem 1.1. Following the notation introduced in Section 4, consider the weighted blow-up of $[B_{+}/\mathbb{G}_{m}]\to\mathcal{W}$ along $\bar{J}=(x_{1}^{1/w_{1}},x_{2}^{1/w_{2}})$. Since $B_{+}\subset B=\operatorname{Spec}(\tilde{R})$ with $\mathcal{O}_{\mathcal{W}}\subset\tilde{R}\subset\mathcal{O}_{\mathcal{W}}[t,t^{-1}]$, the co-action $\mathcal{O}_{\mathcal{W}}\to\mathcal{O}_{\mathcal{W}}\otimes\mathcal{O}_{T}$ naturally lifts to a co-action $\tilde{R}\to\tilde{R}\otimes\mathcal{O}_{T}$ by sending $t\mapsto t\otimes 1$. In this way, the $T$-action on $\mathcal{W}$ lifts to an action of $T^{\prime}=T\times\mathbb{G}_{m}$ on $B_{+}$, inducing a morphism $[B_{+}/T^{\prime}]\to[\mathcal{W}/T]$ fitting in a cartesian diagram:
${{[B_{+}/\mathbb{G}_{m}]}}$${{[B_{+}/T^{\prime}]}}$${\mathcal{W}}$${{[\mathcal{W}/T]}}$
Since $\mathcal{W}\to[\mathcal{W}/T]$ is a smooth cover, and $[B_{+}/\mathbb{G}_{m}]\to\mathcal{W}$ is proper and birational by Lemma 4.4, $[B_{+}/T^{\prime}]\to[\mathcal{W}/T]$ is also proper and birational. In order to show $[B_{+}/T^{\prime}]$ has finite stabilizers, it suffices to show its relative inertia over $[\mathcal{W}/T]$ is finite, since by hypothesis $[\mathcal{W}/T]$ has finite stabilizers. Indeed, that is the case, as the relative inertia of $[B_{+}/T^{\prime}]$ over $[\mathcal{W}/T]$ is finite after base change to the smooth cover $\mathcal{W}\to[\mathcal{W}/T]$.
Finally, by part (3) of Theorem 1.3, the order of $\mathcal{Z}^{\prime}\subset\mathcal{W}^{\prime}:=B_{+}$ at every point lying over $q$ is strictly smaller than $a_{1}$. Since every point lying over $\mathcal{Q}$ is in the same $T^{\prime}$-orbit as points lying over $q$, we deduce the same conclusion for points in $\mathcal{Z}^{\prime}$ lying over $\mathcal{Q}$. ∎